Allan wants to randomly choose a pair of pants and a shirt from his closet. The list below shows all possible combinations of pants and shirts, where the first color is for the pants and the second color is for the shirt.


(blue, black)

(blue, blue)

(blue, green)

(blue, green)

(black, black)

(black, blue)

(black, green)

(black, green)


If Allan randomly chooses a pair of pants and a shirt, what is the probability that Allan will choose a black pair of pants and a blue shirt?


Identify a pants-and-shirt combination that has a 1

4


14 probability of being randomly chosen.


1
Expert's answer
2021-04-15T06:56:48-0400

1.


"P(black , blue)=\dfrac{1}{8}"

2.


"\dfrac{1}{4}=\dfrac{2}{8}=P(blue,green)=P(black,green)"

Allan randomly chooses a pair of blue pants and a green shirt.


Or


Allan randomly chooses a pair of black pants and a green shirt.